OpenStudy (anonymous):

Find the polynomial f(x) that has the roots of -2, 3 of multiplicity 2. Explain how you would verify the zeros of f(x).

OpenStudy (anonymous):

(x+2)(x-3)=0 x^2-x-6=0 f(x)=x^2-x-6